A forensic psychologist tests the extent to which the age of a criminal (X) predicts the age of the victim () for nonviolent crimes. The psychologist uses the case files to record the age of five criminals and the age of the victim in those cases. The hypothetical data are listed in the following table Age of Criminal Age of Victim 25 20 25 25 29 16 (a) Compute the method of least squares to find the equation of the regression line. (Round your answers to three decimal places.) (b) Use the regression equation to determine the predicted age of a victim of a nonviolent crime when the criminal is 22 years old. (Round your answer to three decimal places.) years old
Answer #1

x: Age of criminal: Independent variable

y: Age of victim: Dependent variable

a) The regression equation is y = b*x + a

Here a is intercept and b is a slope.

b=-.ry _ nzy 2

x y xy x^2
33 25 825 1089
25 20 500 625
29 25 725 841
18 22 396 324
13 16 208 169
Total 118 108 2654 3048

ry = 2654

sum x^{2} = 3048

ar{x} = 23.6

ar{y} = 21.6

n = 5

2654 -5 * 23.6 21.6 105.2 263.2 0.400 3048-523.62

a-y _ b * 21.6-0.400 * 23.6-12.167

So the regression equation is

hat{y} = 0.400*X + 12.167

b)

Criminal = X = 22

hat{y} = 0.400*X + 12.167 = 0.400*22 + 12.167 = 20.960